a) Suppose the market demand curve for a product is given by 𝑄



𝑑 = 1000 βˆ’ 10𝑃 and the market supply curve is given by 𝑄𝑠 = βˆ’50 + 25𝑃.



i. What are the equilibrium price and quantity [3Marks]



ii. What is the inverse form of the demand curve [2 Marks]




1
Expert's answer
2022-01-24T12:56:43-0500

i. The equilibrium price and quantity are:

Qd = Qs,

1000 βˆ’ 10P = βˆ’50 + 25P,

35P = 150,

P = 4.29,

Q = 1000 - 10Γ—4.29 = 957.14 units.


ii. The inverse form of the demand curve is:

P = 100 - 0.1Q.